# Fermion liquids as quantum Hall liquids in phase space: A unified approach for anomalies and responses

### v1

@@ Abstract
The discovery of many strongly correlated metallic phases have inspired novel routes to generalize or go beyond the celebrated Landau Fermi liquid theory. To this end, from universal consideration of symmetries and anomalies, Else, Thorngren and Senthil (ETS) ETS have introduced a class of theories called ersatz Fermi liquids which possess a Fermi surface and satisfy a generalized Luttinger’s theorem. In this work, we view all such fermion liquids obeying the Luttinger theorem as incompressible quantum Hall liquids in higher-dimensional phase space and use it as the starting point to derive their effective low-energy field theory. The non-commutativity of phase space motivates us to use Seiberg-Witten map to derive the field theory in an ordinary (commutative) space and naturally leads to terms that correspond to the correct topological Chern-Simons action postulated by ETS in one, two, and three dimensions. Additionally, our approach also reproduces all the non-topological terms that characterize important contributions to the response, including the semiclassical equations of motion. Finally, our derivations of Chern-Simons terms from the Seiberg-Witten map also verify a longstanding conjecture in non-commutative field theory.
@@ I Introduction
Metallic phases formed by interacting electrons are of central importance in condensed matter physics and materials science. The most familiar of them is the Landau Fermi liquid, which can be connected smoothly to the free Fermi gas where quasi-electron and hole excitations are well-defined around the Fermi surface Coleman-book ; Yang-book . Recently, many experiments have shed light on a variaty of “non-Fermi liquid” phases where the effects of interactions are nonperturbative, and quasiparticle excitations are generally absent—distinct from the behavior observed in Fermi liquids Lee-NF-review . Nevertheless, many known examples of non-Fermi liquids support Fermi surfaces that satisfy the Luttinger theorem, like the 1D Luttinger liquid Voit-1995 . Due to the nonperturbative nature of interaction, we do not have a general way to study non-Fermi liquids, except in one dimension where the powerful method of bosonization is available Stone-book ; Tsvelik-book ; Giamarchi-book ; vonDelft-review .
@@
In this paper, we demonstrate that considerable insights into fermion liquids (including both Fermi and non-Fermi liquids in any dimensions) can be gained by viewing them as incompressible quantum Hall liquids in phase space, as long as the Luttinger theorem is obeyed. A quantum Hall liquid could not be further from a Fermi liquid at first glance: The former is incompressible with a gap for bulk excitations, whereas the latter is compressible and supports gapless quasiparticle excitations around its Fermi surface in the bulk. The former requires breaking of time-reversal symmetry to exist in two dimensions, while there is no such requirement for the latter. Here, we point out that they are very closely related in the following way: A $d$-dimensional Fermi liquid is an incompressible liquid in phase space, with the Fermi surface being part of its phase space boundary. We argue that this phase space incompressible liquid is nothing but an integer quantum Hall liquid in $2d$ spatial dimensions, whose gapless boundary excitations correspond to the quasiparticle excitations around the Fermi surface of the Fermi liquid. The key connection is the non-commutativity of phase space. We further argue that this analogy can be extended to a class of non-Fermi liquid metallic states termed ersatz Fermi liquid (EFL) by Else, Thorngren, and Senthil (ETS) ETS . There are no sharp quasiparticles there but the Fermi surface remains well defined and satisfies the Luttinger theorem (or more precisely, the Luttinger sum rule) Luttinger-Ward ; Luttinger ; AGD-book . In our picture, the Luttinger theorem is essentially a trivial consequence of the incompressibility in phase space.
@@
It is important to emphasize that our approach here is in the opposite direction to existing literature Qi2015 ; Else-PRL2021 ; Wang-PRB2021 ; Lu2023 ; Yuxuan2024 . Instead of assuming the phase-space Chern-Simons (CS) term and explore its consequences, we take the Luttinger theorem of the fermion liquid as the starting point. Our proposed effective noncommutative field theory for the incompressible liquid in phase space enables us to derive the anticipated CS term that gives rises the ETS anomaly ETS  [^1] A derivation for a very special case can be found in the very short note Wang-arxiv .. More importantly, our derivation yields all other non-topological responses in a controlled and systematic way. Thus, our work complements and significantly extends recent work on non-Fermi liquids through symmetries and anomalies.
@@ I.1 Motivation and nontechnical discussion
Before diving into technicality, we first illustrate the ideas outlined above by considering the simplest case $d=1$ (1D), where the ersatz Fermi liquid is simply the familiar Luttinger liquid. We show that it can be viewed as an integer quantum Hall liquid in $d=2$ (2D), which is the dimension of the phase space of the Luttinger liquid. As shown in Ref. Yang-PRA2020 , we can view the 2D phase space as the 2D real space subject to a strong perpendicular magnetic field, where all particles are confined to the lowest Landau level. In this representation, the two components of the guiding center, ${\bf R}=(R_{x},R_{y})$ represent $(x,p)$ of the original 1D system if we set $\hbar$ and the magnetic length $\ell_{B}$ to be one, as $[R_{x},R_{y}]=i\ell^{2}_{B}$ Yang-book and $[x,p]=i\hbar$. The kinetic energy $p^{2}\sim R_{y}^{2}$ thus becomes a confining potential along the $y$-direction, and the lowest Landau level electrons form a strip of $\nu=1$ quantum Hall liquid with width $2k_{F}$ aligned along the $x$-direction. Note that the Fermi points ±$\pm k_{F}$ become the $y$ coordinates of the edge locations, and the low-energy excitations of the Luttinger liquid are nothing but the familiar (chiral) edge excitations of the quantum Hall liquid, with left and right movers localized in the lower and upper edges respectively. The above discussion is illustrated in Fig. 1. Naively, the left and right moving electrons are separately conserved corresponding to an emergent U(1)$\times$×U(1) symmetry. However, the extra U(1) symmetry (in addition to overall charge conservation) suffers a ’t Hooft anomaly, which, in the 2D formulation is nothing but the bulk-edge correspondence of the quantum Hall liquid: electrons move from the lower edge to the upper edge through the bulk when an electric field is applied in the $x$-direction due to the quantized Hall effect, ruining the separate conservation of the left and right movers. Just like for $2$D $\nu=1$ quantum Hall liquid, this anomaly is captured by the Chern-Simons action of the external gauge potential $A$ in the $2+1$-D phase space for Luttinger liquid:
$S[A]=\frac{1}{4\pi}\int~{}d^{3}x~{}A\wedge dA=\frac{1}{4\pi}\int~{}d^{3}x~{}\epsilon^{\alpha\beta\gamma}A_{\alpha}\partial_{\beta}A_{\gamma}$ (1)
@@
Figure 1: 1D Luttinger liquid (a) as an integer quantum Hall strip (b) in 2D. The strip is aligned along the horizontal (i.e., $x$) direction. The vertical direction is the linear momentum labelled by $k$, and the kinetic energy maps onto a confining potential that restricts the electrons to $-k_{F}<k<k_{F}$ (dashed lines) in the ground state. The low-energy excitations are the edge fluctuations (solid lines) that move to the left and right at the lower and upper edges respectively. There is an anomaly when an electric field $\mathbf{E}$ (in the $x$ direction) is present, which pumps electrons from the lower edge to the upper edge via the bulk quantized Hall conductance in this analogy.
@@
A Fermi liquid or ersatz Fermi liquid that satisfies the Luttinger theorem is, by definition, an incompressible liquid in phase space. This is because the Luttinger theorem states that the momentum space volume enclosed by the Fermi surface is Luttinger
@@
where $n_{r}$ is the real space density of the liquid, implying the phase space density of the liquid
@@
is fixed, as long as we identify the Fermi surface as its boundary in momentum space, as we did in the 1D Luttinger liquid above. Similarly, the kinetic energy (which can correspond to some band dispersion, not necessarily quadratic) in the phase space, that depends on the momentum coordinates only, maps onto a confinement potential in momentum space. As we will demonstrate, the phase space action for such an incompressible liquid includess, among other terms, the topological CS term in $12d+$ 1 spacetime dimensions:
$S_{0}[A]=\frac{i}{(d+1)!(2\pi)^{d}}\int d^{2d+1}x\ A\wedge dA\cdots\wedge dA,$ (5)
@@
At this point, let us briefly review the core result of Ref. ETS and related work. ETS identify that the microscopic translation symmetry in an ersatz Fermi liquid maps to a global symmetry $G_{\mathrm{IR}}$ of the infrared (IR) theory. In two spatial dimensions, the Fermi surface is parameterized by a single parameter $\theta$, and the appropriate symmetry is the loop group LU(1) of U(1). LU(1) is the set of all smooth maps from the circle $S^{1}$ to U(1), and reflects the fact that the quasiparticle density is conserved at each point on the Fermi surface. The kinematic properties of the ersatz Fermi liquids are governed by the ‘t Hooft anomaly of this emergent loop group symmetry, and the CS term in $12d+$ 1 space-time dimensions was postulated as the appropriate topological term describing the higher-dimensional symmetry-protected topological (SPT) phase whose boundary theory carries the ’t Hooft anomaly for the $G_{\mathrm{IR}}$ gauge fields. The interpretation of the Chern-Simons action as a phase-space quantum Hall liquid was mentioned in passing in Ref. ETS and in earlier literature. Specifically, Ref. Qi2015 analyzed the responses arising from the phase space Chern Simons term extensively. However, neither of these works take into consideration the non-commutativity of phase space. Using the analogy with (higher-dimensional) quantum Hall effect and the inherent non-commutativity of phase space, we are able to derive Eq. (5). This is the primary goal and main result in this work.
@@
The organization of this manuscript is as follows. In Sec. II, we provide a brief review of noncommutative field theory and the Seiberg-Witten (SW) map Seiberg-Witten . This serves as a comprehensive introduction to the key technique that we will use in formulating the field theory for the fermion liquid in noncommutative phase space. Readers who are familiar with noncommutative field theory may skip this section. In Sec. III, we introduce the phase-space action in arbitrary dimensions, and discuss general considerations that hold for all the special cases discussed in later sections. In Sec. IV, we demonstrate the power of our approach by formulating a 1D fermion liquid (i.e., a Luttinger liquid) as the corresponding noncommutative field theory in two-dimensional phase space. By assuming that the fermion liquid satisfies Luttinger’s theorem, we show explicitly how the CS term emerges naturally from the SW map. In this sense, we successfully derive the ETS anomaly in one-dimensional ersatz Fermi liquid ETS . Then, the same approach is employed for two-dimensional fermion liquids in Sec. V. Here, we obtain the most important result: The CS term governing the anomaly and universal topological responses emerges from the second-order SW map. Along the way, we also obtain other terms that describe non-topological responses. These terms describe the usual electromagnetic responses in the semiclassical theory of metals, which are relevant and important in realistic systems. Next, we discuss fermion liquids in 3D and beyond in Sec. VI. Parallel to the 2D case, we obtain the 7D CS term from the third-order expansion of the SW map. Lastly, we summarize our work in Sec. VII. The Appendices discuss some technical issues and mathematical derivations using the SW map, which also form an important part of the present work.
@@ II Review of noncommutative field theory
A theory defined in non-commutative space $X^{\mu}$ can be traded for a non-commutative field theory defined on the usual commuting coordinates $x^{\mu}$ ($[x^{\mu},x^{\nu}]=0$). For this, the product between two noncommutative fields needs to be replaced by the Weyl-Moyal star product, which is defined as
@@ II.1.1 Gauge equivalence conditions
In order to map or approximate a noncommutative field theory by a commutative field theory, Eqs. (15), (II.1.1) and (II.1.1) need to be solved consistently. This is achieved by assuming the following functional forms for the noncommutative fields and gauge parameters Seiberg-Witten :
@@
Then, the noncommutative fields are expanded order-by-order in $\theta$ such that one recovers the commutative fields as $\theta\rightarrow 0$:
@@
The symbol $\Lambda^{(k)}_{\lambda}$ denotes a term in the $k$-th order of $\theta$. For $k=0$, we have $\Lambda_{\lambda}^{(0)}=\lambda$. For $k>0$, $\Lambda^{(k)}_{\lambda}$ is a function of $\lambda$, $A_{\mu}$, and their temporal and spatial derivatives. In principle, Eqs. (15), (II.1.1) and (II.1.1) can then be solved order by order in $\theta$. Nevertheless, the task turns out to be very challenging when one considers higher order terms. In Appendix A, we discuss some technical and mathematical results with respect to the SW map, which are important to our work.

### v2

@@ Abstract
The discovery of many strongly correlated metallic phases has inspired different routes to generalize or go beyond the celebrated Landau Fermi liquid theory. To this end, from universal consideration of symmetries and anomalies, Else, Thorngren and Senthil (ETS) have introduced a class of theories called ersatz Fermi liquids which possess a Fermi surface and satisfy a generalized Luttinger’s theorem. In this work, we view all such fermion liquids obeying the Luttinger theorem as incompressible quantum Hall liquids in higher-dimensional phase space and use it as the starting point to derive their effective low-energy field theory. The noncommutativity of phase space motivates us to use the Seiberg-Witten map to derive the field theory in an ordinary (commutative) space and naturally leads to terms that correspond to the correct topological Chern-Simons action postulated by ETS in one, two, and three dimensions. Additionally, our approach also reproduces all the non-topological terms that characterize important contributions to the response, including the semiclassical equations of motion. Finally, our derivations of Chern-Simons terms from the Seiberg-Witten map also verify a longstanding conjecture in noncommutative field theory.
@@ I Introduction
Metallic phases formed by interacting electrons are of central importance in condensed matter physics and materials science. The most familiar of them is the Landau Fermi liquid, which can be connected smoothly to the free Fermi gas where quasi-electron and hole excitations are well defined around the Fermi surface Coleman-book ; Yang-book . Recently, many experiments have shed light on a variety of “non-Fermi liquid” phases where the effects of interactions are nonperturbative, and quasiparticle excitations are generally absent—distinct from the behavior observed in Fermi liquids Lee-NF-review . Nevertheless, many known examples of non-Fermi liquids support Fermi surfaces that satisfy the Luttinger theorem, like the one-dimensional(1D) Luttinger liquid Voit-1995 . Due to the nonperturbative nature of interaction, we do not have a general way to study non-Fermi liquids, except in one dimension where the powerful method of bosonization is available Stone-book ; Tsvelik-book ; Giamarchi-book ; vonDelft-review .
@@
In this paper, we demonstrate that considerable insights into fermion liquids (including both Fermi and non-Fermi liquids in any dimensions) can be gained by viewing them as incompressible quantum Hall liquids in phase space, as long as the Luttinger theorem is obeyed. A quantum Hall liquid could not be further from a Fermi liquid at first glance: The former is incompressible with a gap for bulk excitations, whereas the latter is compressible and supports gapless quasiparticle excitations around its Fermi surface in the bulk. The former requires breaking of time-reversal symmetry to exist in two dimensions, while there is no such requirement for the latter. Here, we point out that they are very closely related in the following way: A $d$-dimensional Fermi liquid is an incompressible liquid in phase space, with the Fermi surface being part of its phase-space boundary. We argue that this phase space incompressible liquid is nothing but an integer quantum Hall liquid in $2d$ spatial dimensions, whose gapless boundary excitations correspond to the quasiparticle excitations around the Fermi surface of the Fermi liquid. The key connection is the noncommutativity of phase space. We further argue that this analogy can be extended to a class of non-Fermi liquid metallic states termed ersatz Fermi liquid (EFL) by Else, Thorngren, and Senthil (ETS) ETS . There are no sharp quasiparticles there but the Fermi surface remains well defined and satisfies the Luttinger theorem (or more precisely, the Luttinger sum rule) Luttinger-Ward ; Luttinger ; AGD-book . In our picture, the Luttinger theorem is essentially a trivial consequence of the incompressibility in phase space.
@@
It is important to emphasize that our approach here is complementary to existing literature on the subject Qi2015 ; Else-PRL2021 ; Wang-PRB2021 ; Lu2023 ; Yuxuan2024 . Instead of motivating the phase-space Chern-Simons (CS) term on the basis of an anomaly-matching argument, our proposed effective noncommutative field theory for the incompressible liquid in phase space enables us to derive the anticipated CS term that gives rise to the ETS anomaly ETS  [^1] A derivation for a very special case can be found in the very short note Wang-arxiv .. More importantly, our derivation yields all other non-topological responses in a controlled and systematic way. Thus, our work complements and significantly extends recent work on non-Fermi liquids through symmetries and anomalies.
@@ I.1 Motivation and nontechnical discussion
Before diving into technicality, we first illustrate the ideas outlined above by considering the simplest case $d=1$ (1D), where the ersatz Fermi liquid is simply the familiar Luttinger liquid. We show that it can be viewed as an integer quantum Hall liquid in $d=2$ (2D), which is the dimension of the phase space of the Luttinger liquid. As shown in Ref. Yang-PRA2020 , we can view the 2D phase space as the 2D real space subject to a strong perpendicular magnetic field, where all particles are confined to the lowest Landau level. In this representation, the two components of the guiding center, ${\bf R}=(R_{x},R_{y})$ represent $(x,p)$ of the original 1D system if we set $\hbar$ and the magnetic length $\ell_{B}$ to be one, as $[R_{x},R_{y}]=i\ell^{2}_{B}$ Yang-book and $[x,p]=i\hbar$. The kinetic energy $p^{2}\sim R_{y}^{2}$ thus becomes a confining potential along the $y$ direction, and the lowest Landau level electrons form a strip of $\nu=1$ quantum Hall liquid with width $2k_{F}$ aligned along the $x$ direction. Note that the Fermi points ±$\pm k_{F}$ become the $y$ coordinates of the edge locations, and the low-energy excitations of the Luttinger liquid are nothing but the familiar (chiral) edge excitations of the quantum Hall liquid, with left and right movers localized in the lower and upper edges respectively. The above discussion is illustrated in Fig. 1. Naively, the left and right moving electrons are separately conserved corresponding to an emergent U(1)$\times$×U(1) symmetry. However, the extra U(1) symmetry (in addition to overall charge conservation) suffers a ’t Hooft anomaly, which, in the 2D formulation is nothing but the bulk-edge correspondence of the quantum Hall liquid: electrons move from the lower edge to the upper edge through the bulk when an electric field is applied in the $x$ direction due to the quantized Hall effect, ruining the separate conservation of the left and right movers. Just like for $2$D $\nu=1$ quantum Hall liquid, this anomaly is captured by the Chern-Simons action of the external gauge potential $A$ in the $2+1$-D phase space for Luttinger liquid:
$S[A]=\frac{i}{4\pi}\int~{}A\wedge dA=\frac{i}{4\pi}\int~{}d^{3}x~{}\epsilon^{\alpha\beta\gamma}A_{\alpha}\partial_{\beta}A_{\gamma}$ (1)
@@
Figure 1: One-dimensional Luttinger liquid (a) as an integer quantum Hall strip (b) in 2D. The strip is aligned along the horizontal (i.e., $x$) direction. The vertical direction is the linear momentum labeled by $k$, and the kinetic energy maps onto a confining potential that restricts the electrons to $-k_{F}<k<k_{F}$ (dashed lines) in the ground state. The low-energy excitations are the edge fluctuations (solid lines) that move to the left and right at the lower and upper edges, respectively. There is an anomaly when an electric field $\mathbf{E}$ (in the $x$ direction) is present, which pumps electrons from the lower edge to the upper edge via the bulk quantized Hall conductance in this analogy.
@@
A Fermi liquid or ersatz Fermi liquid that satisfies the Luttinger theorem is, by definition, an incompressible liquid in phase space. This is because the Luttinger theorem states that the momentum-space volume enclosed by the Fermi surface is Luttinger
@@
where $n_{r}$ is the real-space density of the liquid, implying the phase-space density of the liquid
@@
is fixed, as long as we identify the Fermi surface as its boundary in momentum space, as we did in the 1D Luttinger liquid above. Similarly, the kinetic energy (which can correspond to some band dispersion, not necessarily quadratic) in the phase-space, that depends on the momentum coordinates only, maps onto a confinement potential in momentum space. As we will demonstrate, the phase-space action for such an incompressible liquid includes, among other terms, the topological CS term in $12d+$ 1 spacetime dimensions:
$S_{0}[A]=\frac{i}{(d+1)!(2\pi)^{d}}\int A\wedge dA\cdots\wedge dA,$ (5)
@@
At this point, let us briefly review the core result of Ref. ETS and related work. ETS identify that the microscopic translation symmetry in an ersatz Fermi liquid maps to a global symmetry $G_{\mathrm{IR}}$ of the infrared (IR) theory. In two spatial dimensions, the Fermi surface is parametrized by a single parameter $\theta$, and the appropriate symmetry is the loop group LU(1) of U(1). LU(1) is the set of all smooth maps from the circle $S^{1}$ to U(1), and reflects the fact that the quasiparticle density is conserved at each point on the Fermi surface. The kinematic properties of the ersatz Fermi liquids are governed by the ‘t Hooft anomaly of this emergent loop group symmetry, and the CS term in $12d+$ 1 space-time dimensions was postulated as the appropriate topological term describing the higher-dimensional symmetry-protected topological (SPT) phase whose boundary theory carries the ’t Hooft anomaly for the $G_{\mathrm{IR}}$ gauge fields. The interpretation of the Chern-Simons action as a phase-space quantum Hall liquid was mentioned in passing in Ref. ETS and in earlier literature. Specifically, Ref. Qi2015 analyzed the responses arising from the phase space Chern Simons term extensively. However, neither of these works take into consideration the noncommutativity of phase space. Using the analogy with (higher-dimensional) quantum Hall effect and the inherent noncommutativity of phase space, we are able to derive Eq. (5). This is the primary goal and main result in this work.
The higher-dimensional CS action has been discussed in literature in the context of higher-dimensional generalizations of the quantum Hall effect Zhang-Hu . Earlier works by Karabali and Nair KN01 ; KN02 ; KN03 have argued that the $(2k+1)$-dimensional CS term describing the dynamics of a quantum Hall droplet on complex projective spaces $\textbf{CP}^{k}$ arises from a Seiberg-Witten type mapping. While the analogy to the higher-dimensional quantum Hall effect is important to our work as a motivation, our starting point is much simpler, that of an incompressible liquid in phase space. The key derivations rely heavily on the symplectic form for the noncommutative parameter that characterizes quantum phase space.
@@
The organization of this manuscript is as follows. In Sec. II, we provide a brief review of noncommutative field theory and the Seiberg-Witten (SW) map Seiberg-Witten . This serves as a comprehensive introduction to the key technique that we will use in formulating the field theory for the fermion liquid in noncommutative phase space. Readers who are familiar with noncommutative field theory may skip this section. In Sec. III, we introduce the phase-space action in arbitrary dimensions, and discuss general considerations that hold for all the special cases discussed in later sections. In Sec. IV, we demonstrate the power of our approach by formulating a 1D fermion liquid (i.e., a Luttinger liquid) as the corresponding noncommutative field theory in two-dimensional phase space. By assuming that the fermion liquid satisfies Luttinger’s theorem, we show explicitly how the CS term emerges naturally from the SW map. In this sense, we successfully derive the ETS anomaly in one-dimensional ersatz Fermi liquid ETS . Then, the same approach is employed for two-dimensional fermion liquids in Sec. V. Here, we obtain the most important result: The CS term governing the anomaly and universal topological responses emerges from the second-order SW map. Along the way, we also obtain other terms that describe non-topological responses. These terms describe the usual electromagnetic responses in the semiclassical theory of metals, which are relevant and important in realistic systems. Next, we discuss fermion liquids in three dimensions (3D) and beyond in Sec. VI. Parallel to the 2D case, we obtain the seven-dimensional (7D) CS term from the third-order expansion of the SW map. Lastly, we summarize our work in Sec. VII. The Appendixes discuss some technical issues and mathematical derivations using the SW map, which also form an important part of the present work.
@@ II Review of noncommutative field theory
A theory defined in noncommutative space $X^{\mu}$ can be traded for a noncommutative field theory defined on the usual commuting coordinates $x^{\mu}$ ($[x^{\mu},x^{\nu}]=0$). For this, the product between two noncommutative fields needs to be replaced by the Weyl-Moyal star product, which is defined as
@@ II.1.1 Gauge equivalence conditions
In order to map or approximate a noncommutative field theory by a commutative field theory, Eqs. (15), (II.1.1), and (II.1.1) need to be solved consistently. This is achieved by assuming the following functional forms for the noncommutative fields and gauge parameters Seiberg-Witten :
@@
Then, the noncommutative fields are expanded order by order in $\theta$ such that one recovers the commutative fields as $\theta\rightarrow 0$:
@@
The symbol $\Lambda^{(k)}_{\lambda}$ denotes a term in the $k$-th order of $\theta$. For $k=0$, we have $\Lambda_{\lambda}^{(0)}=\lambda$. For $k>0$, $\Lambda^{(k)}_{\lambda}$ is a function of $\lambda$, $A_{\mu}$, and their temporal and spatial derivatives. In principle, Eqs. (15), (II.1.1), and (II.1.1) can then be solved order by order in $\theta$. Nevertheless, the task turns out to be very challenging when one considers higher order terms. In Appendix A, we discuss some technical and mathematical results with respect to the SW map, which are important to our work.